OpenStudy (jfruscianterhcp):

A garden plot is to contain 240 sq. ft. If its length is to be 3 times its width,what should its dimensions be?

OpenStudy (jfruscianterhcp):

help sir :)

OpenStudy (anonymous):

if you call the width \(w\) then the length would be \(3w\) making the area \[w\times 3w=3w^2\] since you know it is 240 set \[3w^2=240\] and solve for \(w\) om twp steps

OpenStudy (anonymous):

in two steps

OpenStudy (anonymous):

divide by 3 take the square root

OpenStudy (jfruscianterhcp):

it should be two answers

OpenStudy (jfruscianterhcp):

what are the dimensions

OpenStudy (anonymous):

we are solving for one variable, namely \(w\) which is the width

OpenStudy (anonymous):

if you know \(w\) then the length is three time that

OpenStudy (jfruscianterhcp):

the answer should be 4 square root of 5 and 12 square root of 5

OpenStudy (anonymous):

yes it should

OpenStudy (anonymous):

is it clear how to get that answer?

OpenStudy (jfruscianterhcp):

how to get 12 square root of 5?

OpenStudy (anonymous):

multiply \(4\sqrt{5}\) by \(3\)

OpenStudy (anonymous):

the width is \(4\sqrt{5}\) and you are told that the length is three times the width, so if you know the width, multiply it by 3 to get the length

OpenStudy (jfruscianterhcp):

why?

OpenStudy (anonymous):

is it clear why the width is \(4\sqrt 5\) ?

OpenStudy (jfruscianterhcp):

yes,but how did you get 12 square root of 5?

OpenStudy (anonymous):

\[\large \text{" its length is to be 3 times its width"}\]

OpenStudy (anonymous):

so as i wrote above, if you know the width, you know the length. multiply it by three

OpenStudy (jfruscianterhcp):

oh I get it, thank you sir, I thought its complicated

OpenStudy (jfruscianterhcp):

sir I have a question

OpenStudy (jfruscianterhcp):

Determine the diagonal of the rectangle inscribed in an isosceles right triangle,if the upper two vertices of the rectangle lie at the midpoints of the legs of the triangle